If 0.007 g of KHP (molar mass = 204.23) requires 30.0 mL of NaOH upon titration, what is the normality of NaOH?


1
Expert's answer
2021-06-23T05:20:18-0400

Normality (N) = m /V × 1/eq

= 0.007/0.03 × 1/Eq

= 39.997 × 0.2333

= 9.333